OpenStudy (anonymous):

the sum of two angles is 180. one of the angles measure 96 degrees. what is the measure of the second angle?

OpenStudy (anonymous):

84˚ degrees

OpenStudy (anonymous):

96+4=100+80=180 180-96=84

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

no prob can i get a medal?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks